# Universal Maximum Likelihood (List) Decoding via Fast Vector-Matrix Multiplication

### v2

@@ I Introduction
When the noise model is known, the likelihood measure on memoryless channels can be equivalently expressed as a distance metric—for instance, Hamming distance for memoryless symmetric channels with bit-flip noise, and Euclidean distance for additive white Gaussian noise (AWGN) channels. For example, in a binary symmetric channel (BSC), ML decoding reduces to minimum-distance decoding: the most likely codeword is the one that is closest in Hamming distance to the received sequence. For linear codes, a refinement is possible. When the code rate is $R=k/n\leq 1/2$, ESD involves checking all $2^{k}$ codewords. When $R>1/2$, it is more efficient to consider the $2^{n-k}$ possible syndromes and perform syndrome decoding, which is also a minimum distance decoder and thus ML. In summary, the complexity of ESD for binary linear codes is well known to be $\mathcal{O}(2^{\min\{k,\,n-k\}}n)$ [10]. We emphasize that this refinement applies only to linear codes, since syndrome decoding is not defined for nonlinear codes. A key characteristic of the ESD is that it is universal [^1] In Information theory, the term universal decoding typically refers to decoding without explicit knowledge of the channel law [26, 37]. In this work, however, we use universal to mean that the decoder is not tied to any specific codebook, consistent with its usage in, for example, [43]. In any case, the intended meaning should be clear from the context [15]. (i.e., code-agnostic) as it does not exploit any algebraic structure of the code, and thus its performance depends only on the code size rather than its specific construction.
@@ I-A Brief Historical Prospective
Several techniques have been proposed to reduce the complexity of ML decoding in the high-SNR regime, typically at the cost of a slight increase in error probability. Such methods are therefore commonly referred to as approximate ML decoding, or near-optimum decoding. While asymptotically optimal at high SNR, these methods perform worse than exact ML decoding at low SNR or for practical block lengths, and the gap in error performance compared to MLD increases with the code dimension. Nevertheless, the worst-case complexity (e.g., in low-SNR conditions) remains essentially the same as that of ESD; see, for example, [9, 13, 20].
@@ I-B The Proposed Algorithm and its Relation to the Prior Work
Unlike all previously known ML decoding methods, the proposed decoder universally achieves a complexity of $\mathcal{O}(2^{k})$, which is independent of the block length $n$ when the code rate $k/n$ is assumed to be fixed. This complexity is also independent of the SNR level, distinguishing it from approaches like GRAND [14]. Moreover, the algorithm does not depend on the specific received sequence, unlike the algebraic ML decoder of [27]. The complexity of our proposed method is smaller by a logarithmic factor compared to the best known trellis-based ML decoding via the Viterbi algorithm. This makes the method particularly attractive for low-SNR channels—a relatively unexplored regime where reliable transmission requires low-rate codes with large minimum distance. For example, Reed–Muller (RM) codes $\text{RM}(r,m)$ have long been employed in deep-space and satellite communications due to their high minimum distance and simple majority-logic decoding [36, 38]. For fixed order $r$, the dimension of RM codes scales as $k=\Omega(\log n)$, i.e., it grows logarithmically with the block length $n$. In this regime of very low SNRs and code rates, approximate ML decoding algorithms fail to provide satisfactory performance, and exact ML decoders cannot achieve lower complexity than exhaustive search. Moreover, when extended to list decoding, the proposed algorithm’s complexity increases only logarithmically with the list size, unlike other approaches such as SOGRAND [53] and List Viterbi [45], whose complexity grows linearly with the list size.
@@
The proposed method is especially relevant for emerging applications such as the ultra-reliable low-latency communication (URLLC), a foundational requirement of 5G and future 6G networks. These systems face a fundamental tradeoff: to increase reliability, one often lengthens the block length $n$, which, for conventional universal decoders, directly increases decoding latency, often scaling as $\mathcal{O}\bigl(2^{k}n\bigr)$. In contrast, our method’s $\mathcal{O}(2^{k})$ complexity decouples decoding delay from the block length. There are two key advantages: 1) for a given message/information size $k$, designers can increase $n$ to boost reliability without increasing decoding delay; and 2) the $n$-fold reduction in complexity becomes increasingly significant for large $n$. These advantages make the proposed method a promising enabler for applications that demand both high reliability and ultra-low latency, including Remote surgery & Tactile Internet, Augmented/Virtual/Mixed Reality (AR/VR/XR), Vehicle-to-everything (V2X) communication, and Industrial automation, among many others [31, 34, 54].
@@ I-C Organization
The core contributions of this paper are presented in Section II. Part II-A introduces the universal decoding method for discrete memoryless channels and analyzes its complexity. Part II-B extends the framework to soft-decision decoding and list-$\ell$ ML decoding, and derives the associated complexities. Part II-C applies the method to erasure decoding. Part II-D focuses on linear codes over symmetric channels, showing how syndrome decoding can be used to further reduce complexity. Finally, Part II-E treats the generalization to $L$-tap ISI channels. We conclude the paper in Section IV, where we also discuss the inherent difficulty of the underlying vector-matrix multiplication routine and, drawing on prior work—though without formal proof—argue that further substantial improvements in complexity are unlikely. Finally, a brief overview of the Mailman algorithm is given in Appendix The Mailman Algorithm for Fast vector-matrix Multiplication.
@@ II-B Extension to Soft Decoding and List Decoding
Observe that the output alphabet $\mathcal{Y}$ and the (conditional) transition probabilities can be arbitrary. Particularly, if the received vector $\boldsymbol{y}$ is considered prior to demodulation—for instance, as the real-valued output of a matched filter—then the output alphabet becomes continuous, i.e., $\mathcal{Y}=\mathbb{R}$. This scenario corresponds to soft decoding, where the channel transition probability $P_{Y|X}$ becomes a conditional probability density function $p_{Y|X}$. It is well known—see, for example, [41, Ch. 4]—that ML decoding in this setting is
@@ III Simulation
We evaluate the performance of the proposed decoder against exhaustive search and GRAND [14] over a memoryless binary symmetric channel (BSC), where each bit is independently flipped with probability $p$. The BSC is chosen primarily for implementation simplicity. To ensure a fair comparison, no attempt was made to significantly optimize or parallelize any of the decoding algorithms.
@@
  • Exhaustive search: Upon receiving a vector $\boldsymbol{y}$, the decoder scans the entire codebook $\mathscr{C}$ and outputs the codeword that minimizes the Hamming distance to $\boldsymbol{y}$.
  • GRAND: The decoder generates candidate error patterns $\boldsymbol{e}\in\{0,1\}^{n}$ in decreasing order of likelihood. Over a memoryless BSC with $p<\tfrac{1}{2}$, this ordering coincides with increasing Hamming weight, with ties broken arbitrarily. Decoding terminates when $\boldsymbol{y}-\boldsymbol{e}$ is a valid codeword, i.e.,
    $\boldsymbol{H}(\boldsymbol{y}-\boldsymbol{e})^{\top}=\boldsymbol{0}_{n-k},$
    where $\boldsymbol{H}$ is the code’s parity check matrix, or when a predefined maximum number of guesses, denoted by max_guess, is reached.
@@
Comparison with Exhaustive Search.
The first plot considers first-order Reed–Muller codes. Since both exhaustive search and the proposed decoder necessarily evaluate all codewords in the codebook, the code rate does not affect the fairness of this comparison. Moreover, for both methods, the bit-flip probability $p$ does not influence the decoding time, which depends only on the block length $n$. The proposed decoder exhibits a noticeably smaller variance in running time; for a fixed block length, any residual variation is attributable solely to implementation-level effects.
@@
The second plot considers second-order Reed–Muller codes, which yield moderate-to-high code rates ($R\geq\tfrac{1}{2}$), corresponding to the regime in which GRAND is typically most effective. In this setting, the bit-flip probability $p$ must remain relatively small, since highly noisy channels require codes with large minimum distance, which high-rate codes cannot provide as a consequence of the Singleton bound.
For this reason, the largest block length considered is $n=32$. When $n=64$ and $p=0.065$, the expected number of bit flips is approximately $64\times 0.065\approx 4$. In such cases, GRAND must examine roughly
@@
The comparison with GRAND is therefore more nuanced. Recall that GRAND sequentially subtracts error patterns in decreasing order of likelihood until a valid codeword is found. Consequently, GRAND performs best when $p$ is small or when the code rate is high. Our simulations show that for smaller block lengths and $p=0.03$, GRAND achieves performance comparable to that of the proposed method, since only a limited number of low-weight error patterns need to be tested. However, when the block length increases to $n\geq 32$ and $p=0.065$, the number of plausible error patterns grows rapidly, leading to a substantial degradation in GRAND’s performance relative to the proposed decoder.
(a) Our proposed method vs. exhaustive search
@@
Figure 1: Decoding time performance for our proposed algorithm, exhaustive search, and GRAND in decoding the outputs of binary symmetric channels with various crossover probabilities $p$, using first-order Reed-Muller codes $\text{RM}(1,m)$.
@@ IV Conclusion
We developed an ML decoding algorithm for arbitrary block codes with $n$-fold complexity reduction for an length-$n$ code. From a practical perspective, the proposed framework is well-suited to latency-critical applications where fast and deterministic decoding is essential. In particular, ultra-reliable low-latency communication (URLLC) systems—central to 5G and beyond—require sub-millisecond decoding to support diverse delay-sensitive domains such as the Tactile Internet, industrial automation, and vehicle-to-everything (V2X) communication [31, 34, 54]. The proposed approach’s $n$-fold reduction in complexity allows system designers to increase redundancy (and thus reliability) without incurring latency penalties, offering a direct advantage over iterative or trellis-based decoding. The framework’s simplicity and reliance on vector-matrix multiplication also make it amenable to efficient implementation on parallel hardware such as GPUs or FPGAs, enabling practical deployment in real-time communication and storage systems where both reliability and latency are critical.
@@ Acknowledgment
This was supported in part by NSF-BSF grant FET-2120262. We thank Swapnil Saha, Michael Schleppy, Mursalin Habib, and Karthik C. S. for valuable input.

### v3

@@ I Introduction
When the noise model is known, the likelihood measure of memoryless channels equivalently reduces to a distance metric—for instance, Hamming distance for memoryless symmetric channels with bit-flip noise, and Euclidean distance for additive white Gaussian noise (AWGN) channels. For example, in a binary symmetric channel (BSC), ML decoding reduces to minimum-distance decoding: the most likely codeword is the one that is closest in Hamming distance to the received sequence. For linear codes, a refinement is possible. When the code rate is $R=k/n\leq 1/2$, ESD involves checking all $2^{k}$ codewords. When $R>1/2$, it is more efficient to consider the $2^{n-k}$ possible syndromes and perform syndrome decoding, which is also a minimum distance decoder and thus ML. In summary, the complexity of ESD for binary linear codes is well known to be $\mathcal{O}(2^{\min\{k,\,n-k\}}n)$ [10]. We emphasize that this refinement applies only to linear codes, since syndrome decoding is not defined for nonlinear codes. A key characteristic of the ESD is that it is universal [^1] In Information theory, the term universal decoding typically refers to decoding without explicit knowledge of the channel law [28, 40]. In this work, however, we use universal to mean that the decoder is not tied to any specific codebook, consistent with its usage in, for example, [46]. In any case, the intended meaning should be clear from the context [15]. (i.e., code-agnostic) as it does not exploit any algebraic structure of the code, and thus its performance depends only on the code size rather than its specific construction.
@@ I-A Brief Historical Prospective
Several techniques have been proposed to reduce the complexity of ML decoding in the high-SNR regime, typically at the cost of a slight increase in error probability. Such methods are therefore commonly referred to as approximate ML decoding, or near-optimum decoding. While asymptotically optimal at high SNR, these methods perform worse than exact ML decoding at low SNR or for practical block lengths, and the gap in error performance compared to MLD increases with the code dimension. Nevertheless, the worst-case complexity (e.g., in low-SNR conditions) remains essentially the same as that of ESD; see, for example, [9, 13, 20, 50].
@@ I-B The Proposed Algorithm and its Relation to the Prior Work
Unlike all previously known ML decoding methods, the proposed decoder universally achieves a complexity of $\mathcal{O}(2^{k})$, which is independent of the block length $n$ when the code rate $k/n$ is assumed to be fixed. This complexity is also independent of the SNR level, distinguishing it from approaches like GRAND [14]. Moreover, the algorithm does not depend on the specific received sequence, unlike the algebraic ML decoder of [29]. The complexity of our proposed method is smaller by a logarithmic factor compared to the best known trellis-based ML decoding via the Viterbi algorithm. This makes the method particularly attractive for low-SNR channels—a relatively unexplored regime where reliable transmission requires low-rate codes with large minimum distance [27]. For example, Reed–Muller (RM) codes $\text{RM}(r,m)$ have long been employed in deep-space and satellite communications due to their high minimum distance and simple majority-logic decoding [39, 41, 35]. For fixed order $r$, the dimension of RM codes scales as $k=\Omega(\log n)$, i.e., it grows logarithmically with the block length $n$. In this regime of very low SNRs and code rates, approximate ML decoding algorithms fail to provide satisfactory performance, and exact ML decoders cannot achieve lower complexity than exhaustive search. Moreover, when extended to list decoding, the proposed algorithm’s complexity increases only logarithmically with the list size, unlike other approaches such as SOGRAND [58] and List Viterbi [48], whose complexity grows linearly with the list size.
@@
The proposed ML decoding framework is particularly well suited for ultra-reliable low-latency communication (URLLC), a foundational requirement of 5G and emerging 6G systems. In conventional universal decoders, improving reliability by increasing the block length $n$ typically incurs a proportional increase in decoding latency, often scaling as $\mathcal{O}(2^{k}n)$. In contrast, the proposed method achieves an $n$-fold reduction in complexity, yielding $\mathcal{O}(2^{k})$ decoding time and effectively decoupling latency from the block length. This allows system designers to increase redundancy and reliability without incurring latency penalties, offering a clear advantage over iterative or trellis-based decoders. Owing to its deterministic runtime and reliance on structured vector–matrix multiplication, the framework is well suited for efficient implementation on parallel hardware such as GPUs or FPGAs [25]. These features position the proposed decoder as a promising enabler for latency-critical applications that demand both high reliability and ultra-low latency, including the Tactile Internet, remote surgery, AR/VR/XR, vehicle-to-everything (V2X) communication, and industrial automation [33, 37, 59].
@@ I-C Organization
The core contributions of this paper are presented in Section II. Part II-A introduces the universal decoding method for discrete memoryless channels and analyzes its complexity. Part II-B extends the framework to soft-decision decoding and list-$\ell$ ML decoding, and derives the associated complexities. Part II-C applies the method to erasure decoding. Part II-D focuses on linear codes over symmetric channels, showing how syndrome decoding can be used to further reduce complexity. Part II-E then treats the application of the proposed framework to channels with memory, namely, $L$-tap ISI channels. Section III provides simulation results validating the theoretical analysis. We conclude the paper in Section IV, where we also discuss the inherent difficulty of the underlying vector-matrix multiplication routine and, drawing on prior work—though without formal proof—argue that further substantial improvements in complexity are unlikely. Finally, a brief overview of the Mailman algorithm is given in Appendix The Mailman Algorithm for Fast vector-matrix Multiplication.
@@ II-B Extension to Soft Decoding and List Decoding
Observe that the output alphabet $\mathcal{Y}$ and the (conditional) transition probabilities can be arbitrary. Particularly, if the received vector $\boldsymbol{y}$ is considered prior to demodulation—for instance, as the real-valued output of a matched filter—then the output alphabet becomes continuous, i.e., $\mathcal{Y}=\mathbb{R}$. This scenario corresponds to soft decoding, where the channel transition probability $P_{Y|X}$ becomes a conditional probability density function $p_{Y|X}$. It is well known—see, for example, [44, Ch. 4]—that ML decoding in this setting is
@@ III Simulation
We evaluate the performance of the proposed decoder against exhaustive search (ESD) and GRAND [14] over a memoryless binary symmetric channel (BSC), where each bit is independently flipped with probability $p$. The choice of ESD and GRAND is because they are universal decoders, just as the proposed decoder. The BSC can be modeled as $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{e}$, where $\boldsymbol{x},\,\boldsymbol{e}\in\mathbb{F}_{2}^{n}$ are respectively the transmitted codeword and the error pattern, and addition is modulo $2$. To ensure a fair comparison, no attempt was made to significantly optimize or parallelize any of the decoding algorithms.
@@
  • ESD: Upon receiving a vector $\boldsymbol{y}$, the decoder scans the entire codebook $\mathscr{C}$ and outputs the codeword that minimizes the Hamming distance to $\boldsymbol{y}$.
  • GRAND: The decoder generates candidate error patterns $\boldsymbol{e}\in\{0,1\}^{n}$ in decreasing order of likelihood. Over a memoryless BSC with $p<\tfrac{1}{2}$, this ordering coincides with increasing Hamming weight, with ties broken arbitrarily. Decoding terminates when $\boldsymbol{y}-\boldsymbol{e}$ is a valid codeword in codebook $\mathscr{C}$, or when a predefined maximum number of guesses, denoted by max_guess, is reached, in which case a decoding failure is declared (i.e., guessing with abandonment [14, 49]).
@@
Comparison with ESD.
The first plot considers first-order Reed–Muller codes. Since both ESD and the proposed decoder necessarily evaluate all codewords in the codebook, the choice of code rates does not affect the fairness of this comparison. Moreover, for both methods, the bit-flip probability $p$ does not influence the decoding time, which depends only on the codebook size. At each operating point, we report the mean decoding time together with error bars corresponding to one standard deviation. The proposed decoder exhibits a noticeably smaller variance in running time; for a fixed block length, any residual variation is attributable solely to implementation-level effects.
@@
The second plot considers $2^{\text{nd}}$-order RM codes with lengths $n=\{8,16,32\}$, corresponding to moderate-to-high rates ($R\geq\tfrac{1}{2}$), where GRAND is most effective [14]. In this regime, $p$ must remain small, since high-rate codes have lower minimum distance and are not suitable for low-SNR channels [27]. For this reason, the largest block length considered is $n=32$. When $n=64$ and $p=0.07$, the expected number of bit flips is $64\times 0.07>4$. In such cases, GRAND must typically examine at least
@@
The comparison with GRAND is therefore restricted to smaller instances. As shown in Fig. 1(b), for small values of $n$, GRAND achieves performance comparable to the proposed decoder, since only a limited number of low-weight error patterns need to be tested. However, when the block length increases to $n\geq 32$, the rapid growth in the number of plausible error patterns leads to substantially higher decoding time and an increased failure rate for GRAND. Although increasing max_guess can reduce failures, it does so at the cost of considerably higher average decoding time.
(a) Our proposed method vs. ESD
@@
Figure 1: Decoding time performance for our proposed algorithm, ESD, and GRAND in decoding over BSC with various values of bit flip probabilities $p$ and block lengths $n$, using Reed-Muller codes.
@@ IV Conclusion
We develop a maximum-likelihood (ML) decoding algorithm for arbitrary block codes by reformulating the ML decoding task as a vector–matrix multiplication problem, after which identifying the ML codeword reduces to selecting the largest entry of the resulting vector. To accelerate this computation, we employ the Mailman algorithm [34], yielding an $n$-fold reduction in decoding complexity for a length-$n$ code. The proposed framework is particularly well suited for latency-critical applications in modern communication systems. Promising directions for future work include developing approximation methods that reduce the associated storage requirements while ensuring near-optimal performance.
@@ Acknowledgment
This was supported in part by NSF-BSF grant FET-2120262. We thank Swapnil Saha, Mursalin Habib, and Karthik C. S. for valuable input.